# CoD: A Diffusion Foundation Model for Image Compression

### v2

@@ Abstract
Figure 1: Overview of Compression-oriented Diffusion (CoD) foundation models, which are trained from scratch to jointly optimize compression and generation. Rather than a fixed codec, CoD serves as a foundational model for downstream diffusion-based codecs such as DiffC [diffc_original], substantially enhancing their performance by replacing Stable Diffusion.
@@ 1 Introduction
Diffusion image compression [cdc, perco, perco_sd, diffeic, diffc, ddcm, oscar] exploits the strong generative priors of denoising diffusion models [diffusion, score_matching, rf] to achieve realistic image reconstruction. To inherit such generative priors, recent diffusion codecs typically adopt large-scale pretrained text-to-image diffusion models such as Stable Diffusion [SD], demonstrating impressive performance in generative image compression.
However, when using text-to-image diffusion models in compression, the role of text conditions remains unclear. Although early works such as Text+Sketch [text+sketch] and PerCo [perco] claim image captions can serve as a high-level semantic signal to facilitate ultra-low bitrate compression, later studies [onedc, oscar] demonstrate that fine-tuning Stable Diffusion to discard text priors can further improve compression efficiency. Additional evidence comes from the zero-shot diffusion compression framework DiffC [diffc], which theoretically measures the “compression capability” of a denoising diffusion model. When applied to Stable Diffusion, DiffC reveals that text conditions are detrimental to compression performance, especially at low bitrates. These observations suggest that text-conditioned diffusion models are not naturally suitable for compression, and that Stable Diffusion is not the ideal foundation model for diffusion codecs.
@@
where $x$, $y$, and $\hat{x}$ denote the original image, compressed representation, and reconstruction, respectively, and $\Theta$ and $\Phi$ represent the trained parameters of the encoder and decoder. If we set $\Theta$ as an image captioner like BLIP-2 [blip] and $\Phi$ as Stable Diffusion, we note that text-conditioned diffusion can be formulated in the compression formulation, as illustrated in the top left of Figure 1. From the perspective of compression, text conditions introduces two limitations. First, human-generated text is less efficient in describing fine-grained spatial and texture details of natural images. Second, the discrete vocabularies in text is non-differentiable to prevent joint optimization of $\Theta$ and $\Phi$, thus making rate–distortion optimization inapplicable.
@@
Figure 2: Framework overview of CoD in pixel and latent spaces. CoD consists of a condition encoder, an entropy bottleneck, a condition decoder and a diffusion model which is decoupled to DiT backbone and DDT head [ddt]. CoD is trained with rectified flow [rf], where $1-\alpha\%$ samples are trained at $t=0$ to jointly optimize distortion and perception.
Low Training Cost. Compared to Stable Diffusion, CoD is far more efficient to train. Stable Diffusion v1.5 requires about 6,250 A100 GPU days [pixart_alpha], whereas CoD only takes around 20 A100 GPU days (0.3% cost). Unlike text-to-image diffusion that relies on text-image training pairs, CoD is fully self-supervised and only requires image datasets, making it easier to scale. Furthermore, while Stable Diffusion uses proprietary datasets, CoD is trained entirely on open datasets such as ImageNet [imagenet], OpenImages [openimage], and SA-1B [sa1b], making it easy to reproduce and unleashing the potential for future exploration across different diffusion codec architectures and training schemes.
@@
High Compression Performance. As a compression-oriented foundation model, CoD enables high-performance downstream codecs. In this paper, we evaluate it using the state-of-the-art zero-shot diffusion compression method, DiffC [diffc_original, diffc]. Compared to Stable Diffusion, DiffC with CoD achieves significantly better performance, especially at low bitrates. With pixel-space CoD, it even matches VTM [vvc] in PSNR while outperforming previous pixel-space perceptual codecs such as MS-ILLM [illm] in FID. The qualitative results in middle column of Figure 1 demonstrate that pixel-space CoD achieves much higher reconstruction fidelity, highlighting the potential of pixel diffusion to approach the theoretical rate–distortion–perception trade-off.
@@ 2.1 Pipeline
Condition Encoding. An encoder firstly compresses $x$ into compact representations. Following [vqgan], we stack residual blocks [resnet] and attention layers [transformer] to construct the encoder. The output is at $1/32$ resolution of the original images.
Entropy bottleneck. It is designed to maintain an ultra-low bitrate for $y$, encouraging the diffusion model to learn strong generative capabilities. This can be implemented as scalar quantization with entropy models [hyperprior], vector quantization [vqgan], or finite scalar quantization [fsq]. For simplicity, we use vector quantization with a codebook size of $N=2^{4}=16$, corresponding to $4\text{ bits}/(32\times 32)=0.0039\text{ bpp}$.
@@
Diffusion-based reconstruction. The diffusion module denoises a randomly sampled Gaussian noise over $T$ steps under the guidance of condition $c$ to reconstruct $\hat{x}$. Following DDT [ddt], we decouple DiT backbone and DDT head to enhance capacity. DDT performs denoising at $1/16$ resolution, where the condition $c$ is concatenated with the noised input along the channel dimension at each step for guidance.
In the latent space, since VAE latents are already at $1/8$ resolution, we apply a $2\times 2$ patch embedding to downsample them to $1/16$, and upsample the output of DDT head back to $1/8$ before passing to the VAE decoder. In the pixel space, we directly use a $16\times 16$ patch embedding to map the noised images to $1/16$ resolution, and adopt the pixel DDT head from [pixnerd], where each feature predicts a neural field that reconstructs a corresponding $16\times 16$ patch. Since the DiT backbone runs at $1/16$ resolution of original images, pixel-space CoD has a similar computation complexity as in latent space. Details are in the appendix.
@@
Figure 4: Scaling law analysis on the Kodak dataset at a resolution of $256\times 256$. All CoD models are at 0.016 bpp while MS-ILLM is at 0.021 bpp. FID is computed over overlapping $64\times 64$ patches following [ddcm].
@@ 2.2 Diffusion Training
Preliminary. Denoising diffusion models [diffc] model the data distribution by progressively perturbing a clean sample $x$ with Gaussian noise $\varepsilon$, transforming $p(x)$ into a standard Gaussian distribution. At timestep $t$, it is formulated as,
@@
with drift $f(t)$ and diffusion $g(t)$ determined by $\alpha_{t}$ and $\sigma_{t}$. Score-based models [score_matching] learn the score function of $p(x)$ to solve the reverse SDE and generate samples.
Rectified Flow (RF) [rf] reinterprets diffusion as a deterministic transformation. It removes the stochastic term from the SDE and adopts a linear transition schedule:
@@
where $v_{t}^{\mathrm{pred}}$ denotes the predicted velocity field. Theoretically, optimizing $\mathcal{L}_{\mathrm{RF}}$ reduces the Kullback–Leibler (KL) divergence between the reconstructed and original distributions [gentiloni2024theoretical], which can reflect the perception term in rate-distortion-perception (R-D-P) theory [rdp]. Hence, optimizing $\mathcal{L}_{\mathrm{RF}}$ naturally promotes perceptual quality.
@@
Discussion. Unified training is designed for continuous flows, where the probability of sampling a fully noised state $t=0$ approaches zero. We observe that color reconstruction in CoD is largely determined by its first denoising step; thus, the absence of $t=0$ optimization often leads to noticeable color shifts. In contrast, DDPM-based codecs such as PerCo [perco] employ discrete timesteps uniformly sampled as $t_{\mathrm{DDPM}}\in{1,\cdots,1000}$, inherently including optimizing of fully noised state. This can be viewed as a special case of our formulation with $1-\alpha\%=1/1000$.
@@ 2.3 Implementation Details
Datasets. Foundation models require large-scale training to achieve optimal performance. Unlike text-to-image generation that depends on text-image data pairs, CoD is fully self-supervised and requires only image datasets, making large-scale scaling more accessible. In this paper, we use ImageNet-21K [imagenet], OpenImage [openimage] and SA-1B [sa1b], including a total of 22M images. All datasets are publicly available, facilitating reproducibility for future research.
Loss Function. For the information bottleneck, we adopt the codebook commitment loss $\mathcal{L}_{\mathrm{C}}$ [vqgan]. Diffusion training employs the rectified flow loss $\mathcal{L}_{\mathrm{RF}}$ and the representation alignment loss $\mathcal{L}_{\mathrm{REPA}}$ with DINOv2 [dino_v2]. Additionally, an auxiliary head (Figure 2) reconstructs both original pixels and DINOv2 features from the condition $c$ to enhance condition learning. Further analysis of this auxiliary loss $\mathcal{L}_{\mathrm{aux}}$ is in the appendix. The overall objective is
@@
Training Costs. CoD is trained on 4 NVIDIA A100 GPUs for approximately 5 days, totaling around 20 A100 GPU days. This represents only about $0.3\%$ of the training cost of Stable Diffusion v1.5 [SD], primarily because: (1) CoD learns the most informative compression conditions end-to-end, rather than relying on high-dimensional, human-defined text embeddings; (2) the unified rectified flow training stabilizes and accelerates convergence; (3) advanced optimization strategies such as REPA [repa] further improve efficiency.
Diffusion Sampling. For ODE-based sampling, we employ the Adam-like-2nd solver [pixnerd] with 25 steps. We use classifier-free guidance [cfg] with scales of 3.0 for pixel space and 1.25 for latent space to enhance conditional generation.
@@ 3.1 Evaluation
CoD naturally operates as a codec at 0.0039 bpp. We evaluate its compression performance on the Kodak [kodak] dataset at $512\times 512$ resolution. To ensure fair comparison among foundation models, we employ the same SD-VAE and compare our latent-space CoD with Stable-Diffusion-based codecs, including zero-shot DiffC [diffc], DDCM [ddcm], and fine-tuned PerCo (SD) [perco_sd] and OSCAR [oscar]. As shown in Figure 6, CoD achieves superior reconstruction quality. A more detailed quantitative comparison is presented in Section 4.
@@ 3.2 Scaling Law
Most diffusion codecs are constrained by the parameter count of Stable Diffusion (approximately 860M), making it difficult to study parameter scaling. It also remains unclear whether the performance advantages of diffusion codecs over GAN-based codecs (e.g., MS-ILLM [illm]) stem from architectural improvements or simply from model scaling. To explore this, we train pixel-space CoDs with varying channel dimensions at $256\times 256$ resolution for 400k steps and evaluate them on Kodak. As shown in Figure 4, we observe a clear trend: increasing model parameters consistently enhances compression performance. Compared with the GAN-based MS-ILLM (181M parameters), CoD achieves notably better reconstruction quality even with only 49M parameters, confirming that the performance gain primarily originates from algorithmic improvements rather than model size.
@@ 3.4 Revisiting Pixel-Space Diffusion
Recent diffusion codecs typically operate in the latent space of a KL-VAE, achieving impressive results at low bitrates (e.g., <<0.1 bpp). However, their performance is fundamentally constrained by the entropy and reconstruction capacity of the underlying VAE. As shown in Figure 5 (right), DiffC [diffc] built on Stable Diffusion cannot surpass the reconstruction limit of the SD-VAE (PSNR 25.77 dB, LPIPS 0.069), with a bitrate ceiling around 0.26 bpp. Similarly, latent-space CoD remains bounded by VAE quality even up to 0.60 bpp, indicating that latent diffusion codecs inherently operate within a limited quality and bitrate range.
@@ 4 Downstream Compression Performance
In this section, we conduct a comprehensive evaluation of CoD when integrated into downstream compression frameworks. Since most diffusion-based codecs do not release their training code, we employ the zero-shot compression method DiffC [diffc] on CoD for comparison. Additional downstream results are presented in the appendix, e.g., finetuning one-step CoD achieves real-time and SOTA performance (Section B), and CoD-based perceptual loss that significantly enhances MS-ILLM [illm] performance (Section C).
Benchmarks and Baselines. Experiments are conducted on the Kodak [kodak], CLIC2020 test set [clic] and Div2K [div2k]. All images are resized and center-cropped to a resolution of $512\times 512$. We report distortion using PSNR and perceptual quality using LPIPS [lpips], DISTS [dists], and FID [fid]. Following [ddcm], FID is computed on extracted patches ($64\times 64$ for Kodak and $128\times 128$ for CLIC2020). Bitrate is measured in bits per pixel (bpp). In addition to diffusion codecs (DiffC, DDCM, PerCo (SD), and OSCAR; see Section 3.1), we further compare against the latent-space codec DiffEIC [diffeic] and pixel-space codecs including the traditional codec VTM [vvc] and perceptual codecs HiFiC [hific], MS-ILLM [illm], CDC [cdc] and TACO [taco]. We also discuss StableCodec [stablecodec] and OneDC [onedc] which are built on pretrained one-step diffusion models in the appendix.
@@
Surprisingly, our codec achieves comparable BPP–PSNR performance to VTM (BD-Rate is $2.1\%$ using VTM as anchor) while delivering significantly better perceptual quality. In contrast, HiFiC, MS-ILLM and CDC ($\rho=0.9$) improve perceptual quality at the expense of PSNR. TACO removes the GAN [gan] loss to balance distortion and perception, but fails to achieve an optimal trade-off. By comparison, DiffC with pixel-space CoD attains the best balance across all bitrates. Although pixel diffusion has been less explored than latent diffusion, our results demonstrate that CoD highlights the potential of pixel diffusion to achieve a superior rate–distortion–perception trade-off.

### v3

@@ Abstract
Figure 1: Overview of Compression-oriented Diffusion (CoD) foundation models, which are trained from scratch to jointly optimize compression and generation. Rather than a fixed codec, CoD serves as a foundational model for downstream diffusion-based codecs such as DiffC [43], substantially enhancing their performance by replacing Stable Diffusion.
@@ 1 Introduction
Diffusion image compression [54, 7, 26, 32, 46, 37, 15] exploits the strong generative priors of denoising diffusion models [20, 42, 33] to achieve realistic image reconstruction. To inherit such generative priors, recent diffusion codecs typically adopt large-scale pretrained text-to-image diffusion models such as Stable Diffusion [39], demonstrating impressive performance in generative image compression.
However, when using text-to-image diffusion models in compression, the role of text conditions remains unclear. Although early works such as Text+Sketch [29] and PerCo [7] claim image captions can serve as a high-level semantic signal to facilitate ultra-low bitrate compression, later studies [53, 15] demonstrate that fine-tuning Stable Diffusion to discard text priors can further improve compression efficiency. Additional evidence comes from the zero-shot diffusion compression framework DiffC [46], which theoretically measures the “compression capability” of a denoising diffusion model. When applied to Stable Diffusion, DiffC reveals that text conditions are detrimental to compression performance, especially at low bitrates. These observations suggest that text-conditioned diffusion models are not naturally suitable for compression, and that Stable Diffusion is not the ideal foundation model for diffusion codecs.
@@
where $x$, $y$, and $\hat{x}$ denote the original image, compressed representation, and reconstruction, respectively, and $\Theta$ and $\Phi$ represent the trained parameters of the encoder and decoder. If we set $\Theta$ as an image captioner like BLIP-2 [30] and $\Phi$ as Stable Diffusion, we note that text-conditioned diffusion can be formulated in the compression formulation, as illustrated in the top left of Figure 1. From the perspective of compression, text conditions introduces two limitations. First, human-generated text is less efficient in describing fine-grained spatial and texture details of natural images. Second, the discrete vocabularies in text is non-differentiable to prevent joint optimization of $\Theta$ and $\Phi$, thus making rate–distortion optimization inapplicable.
@@
Figure 2: Framework overview of CoD in pixel and latent spaces. CoD consists of a condition encoder, an entropy bottleneck, a condition decoder and a diffusion model which is decoupled to DiT backbone and DDT head [49]. CoD is trained with rectified flow [33], where $1-\alpha\%$ samples are trained at $t=0$ to jointly optimize distortion and perception.
Low Training Cost. Compared to Stable Diffusion, CoD is far more efficient to train. Stable Diffusion v1.5 requires about 6,250 A100 GPU days [8], whereas CoD only takes around 20 A100 GPU days (0.3% cost). Unlike text-to-image diffusion that relies on text-image training pairs, CoD is fully self-supervised and only requires image datasets, making it easier to scale. Furthermore, while Stable Diffusion uses proprietary datasets, CoD is trained entirely on open datasets such as ImageNet [40], OpenImages [27], and SA-1B [25], making it easy to reproduce and unleashing the potential for future exploration across different diffusion codec architectures and training schemes.
@@
High Compression Performance. As a compression-oriented foundation model, CoD enables high-performance downstream codecs. In this paper, we evaluate it using the state-of-the-art zero-shot diffusion compression method, DiffC [43, 46]. Compared to Stable Diffusion, DiffC with CoD achieves significantly better performance, especially at low bitrates. With pixel-space CoD, it even matches VTM [47] in PSNR while outperforming previous pixel-space perceptual codecs such as MS-ILLM [36] in FID. The qualitative results in middle column of Figure 1 demonstrate that pixel-space CoD achieves much higher reconstruction fidelity, highlighting the potential of pixel diffusion to approach the theoretical rate–distortion–perception trade-off.
@@ 2.1 Pipeline
Condition Encoding. An encoder firstly compresses $x$ into compact representations. Following [12], we stack residual blocks [18] and attention layers [45] to construct the encoder. The output is at $1/32$ resolution of the original images.
Entropy bottleneck. It is designed to maintain an ultra-low bitrate for $y$, encouraging the diffusion model to learn strong generative capabilities. This can be implemented as scalar quantization with entropy models [5], vector quantization [12], or finite scalar quantization [34]. For simplicity, we use vector quantization with a codebook size of $N=2^{4}=16$, corresponding to $4\text{ bits}/(32\times 32)=0.0039\text{ bpp}$.
@@
Diffusion-based reconstruction. The diffusion module denoises a randomly sampled Gaussian noise over $T$ steps under the guidance of condition $c$ to reconstruct $\hat{x}$. Following DDT [49], we decouple DiT backbone and DDT head to enhance capacity. DDT performs denoising at $1/16$ resolution, where the condition $c$ is concatenated with the noised input along the channel dimension at each step for guidance.
In the latent space, since VAE latents are already at $1/8$ resolution, we apply a $2\times 2$ patch embedding to downsample them to $1/16$, and upsample the output of DDT head back to $1/8$ before passing to the VAE decoder. In the pixel space, we directly use a $16\times 16$ patch embedding to map the noised images to $1/16$ resolution, and adopt the pixel DDT head from [48], where each feature predicts a neural field that reconstructs a corresponding $16\times 16$ patch. Since the DiT backbone runs at $1/16$ resolution of original images, pixel-space CoD has a similar computation complexity as in latent space. Details are in the appendix.
@@
Figure 4: Scaling law analysis on Kodak at $256\times 256$. All CoD models are at 0.016 bpp while MS-ILLM is at 0.021 bpp.
@@ 2.2 Diffusion Training
Preliminary. Denoising diffusion models [46] model the data distribution by progressively perturbing a clean sample $x$ with Gaussian noise $\varepsilon$, transforming $p(x)$ into a standard Gaussian distribution. At timestep $t$, it is formulated as,
@@
with drift $f(t)$ and diffusion $g(t)$ determined by $\alpha_{t}$ and $\sigma_{t}$. Score-based models [42] learn the score function of $p(x)$ to solve the reverse SDE and generate samples.
Rectified Flow (RF) [33] reinterprets diffusion as a deterministic transformation. It removes the stochastic term from the SDE and adopts a linear transition schedule:
@@
where $v_{t}^{\mathrm{pred}}$ denotes the predicted velocity field. Theoretically, optimizing $\mathcal{L}_{\mathrm{RF}}$ reduces the Kullback–Leibler (KL) divergence between the reconstructed and original distributions [13], which can reflect the perception term in rate-distortion-perception (R-D-P) theory [6]. Hence, optimizing $\mathcal{L}_{\mathrm{RF}}$ naturally promotes perceptual quality.
@@
Discussion. Unified training is designed for continuous flows, where the probability of sampling a fully noised state $t=0$ approaches zero. We observe that color reconstruction in CoD is largely determined by its first denoising step; thus, the absence of $t=0$ optimization often leads to noticeable color shifts. In contrast, DDPM-based codecs such as PerCo [7] employ discrete timesteps uniformly sampled as $t_{\mathrm{DDPM}}\in{1,\cdots,1000}$, inherently including optimizing of fully noised state. This can be viewed as a special case of our formulation with $1-\alpha\%=1/1000$.
@@ 2.3 Implementation Details
Datasets. Foundation models require large-scale training to achieve optimal performance. Unlike text-to-image generation that depends on text-image data pairs, CoD is fully self-supervised and requires only image datasets, making large-scale scaling more accessible. In this paper, we use ImageNet-21K [40], OpenImage [27] and SA-1B [25], including a total of 22M images. All datasets are publicly available, facilitating reproducibility for future research.
Loss Function. For the information bottleneck, we adopt the codebook commitment loss $\mathcal{L}_{\mathrm{C}}$ [12]. Diffusion training employs the rectified flow loss $\mathcal{L}_{\mathrm{RF}}$ and the representation alignment loss $\mathcal{L}_{\mathrm{REPA}}$ with DINOv2 [38]. Additionally, an auxiliary head (Figure 2) reconstructs both original pixels and DINOv2 features from the condition $c$ to enhance condition learning. Further analysis of this auxiliary loss $\mathcal{L}_{\mathrm{aux}}$ is in the appendix. The overall objective is
@@
Training Costs. CoD is trained on 4 NVIDIA A100 GPUs for approximately 5 days, totaling around 20 A100 GPU days. This represents only about $0.3\%$ of the training cost of Stable Diffusion v1.5 [39], primarily because: (1) CoD learns the most informative compression conditions end-to-end, rather than relying on high-dimensional, human-defined text embeddings; (2) the unified rectified flow training stabilizes and accelerates convergence; (3) advanced optimization strategies such as REPA [56] further improve efficiency.
Diffusion Sampling. For ODE-based sampling, we employ the Adam-like-2nd solver [48] with 25 steps. We use classifier-free guidance [21] with scales of 3.0 for pixel space and 1.25 for latent space to enhance conditional generation.
@@ 3.1 Evaluation
CoD naturally operates as a codec at 0.0039 bpp. We evaluate its compression performance on the Kodak [10] dataset at $512\times 512$ resolution. To ensure fair comparison among foundation models, we employ the same SD-VAE and compare our latent-space CoD with Stable-Diffusion-based codecs, including zero-shot DiffC [46], DDCM [37], and fine-tuned PerCo (SD) [26] and OSCAR [15]. As shown in Figure 6, CoD achieves superior reconstruction quality. A more detailed quantitative comparison is presented in Section 4.
@@ 3.2 Scaling Law
Most diffusion codecs are constrained by the parameter count of Stable Diffusion (approximately 860M), making it difficult to study parameter scaling. It also remains unclear whether the performance advantages of diffusion codecs over GAN-based codecs (e.g., MS-ILLM [36]) stem from architectural improvements or simply from model scaling. To explore this, we train pixel-space CoDs with varying channel dimensions at $256\times 256$ resolution for 400k steps and evaluate them on Kodak. As shown in Figure 4, we observe a clear trend: increasing model parameters consistently enhances compression performance. Compared with the GAN-based MS-ILLM (181M parameters), CoD achieves notably better reconstruction quality even with only 49M parameters, confirming that the performance gain primarily originates from algorithmic improvements rather than model size.
@@ 3.4 Revisiting Pixel-Space Diffusion
Recent diffusion codecs typically operate in the latent space of a KL-VAE, achieving impressive results at low bitrates (e.g., <<0.1 bpp). However, their performance is fundamentally constrained by the entropy and reconstruction capacity of the underlying VAE. As shown in Figure 5 (right), DiffC [46] built on Stable Diffusion cannot surpass the reconstruction limit of the SD-VAE (PSNR 25.77 dB, LPIPS 0.069), with a bitrate ceiling around 0.26 bpp. Similarly, latent-space CoD remains bounded by VAE quality even up to 0.60 bpp, indicating that latent diffusion codecs inherently operate within a limited quality and bitrate range.
@@ 4 Downstream Compression Performance
In this section, we conduct a comprehensive evaluation of CoD when integrated into downstream compression frameworks. Since most diffusion-based codecs do not release their training code, we employ the zero-shot compression method DiffC [46] on CoD for comparison. Additional downstream results are presented in the appendix, e.g., finetuning one-step CoD achieves real-time and SOTA performance (Section B), and CoD-based perceptual loss that significantly enhances MS-ILLM [36] performance (Section C).
Benchmarks and Baselines. Experiments are conducted on the Kodak [10], CLIC2020 test set [44] and Div2K [2]. All images are resized and center-cropped to a resolution of $512\times 512$. We report distortion using PSNR and perceptual quality using LPIPS [57], DISTS [9], and FID [19]. Following [37], FID is computed on extracted patches ($64\times 64$ for Kodak and $128\times 128$ for CLIC2020). Bitrate is measured in bits per pixel (bpp). In addition to diffusion codecs (DiffC, DDCM, PerCo (SD), and OSCAR; see Section 3.1), we further compare against the latent-space codec DiffEIC [32] and pixel-space codecs including the traditional codec VTM [47] and perceptual codecs HiFiC [35], MS-ILLM [36], CDC [54] and TACO [28]. We also discuss StableCodec [58] and OneDC [53] which are built on pretrained one-step diffusion models in the appendix.
@@
Surprisingly, our codec achieves comparable BPP–PSNR performance to VTM (BD-Rate is $2.1\%$ using VTM as anchor) while delivering significantly better perceptual quality. In contrast, HiFiC, MS-ILLM and CDC ($\rho=0.9$) improve perceptual quality at the expense of PSNR. TACO removes the GAN [14] loss to balance distortion and perception, but fails to achieve an optimal trade-off. By comparison, DiffC with pixel-space CoD attains the best balance across all bitrates. Although pixel diffusion has been less explored than latent diffusion, our results demonstrate that CoD highlights the potential of pixel diffusion to achieve a superior rate–distortion–perception trade-off.